# Scalable Kernel Logistic Regression with Nyström Approximation: Theoretical Analysis and Application to Discrete Choice Modelling

### v1

@@ 6.4 Comparative of optimisation techniques for Nyström KLR
An important consideration of this study is that to obtain the graphs presented, it was necessary to report the value of the objective function at every iteration of the estimation methods. However, obtaining the value of the objective function using the KLR method with large datasets is computationally expensive. Consequently, all estimation methods display CPU times that surpass those that would have been the case if the objective function’s value were not reported, or if it were computed only at intervals of $\Delta_{t}$ iterations.
@@ 7 Conclusion and future work
Firstly, a theoretical study was conducted to characterise the solution set of the KLR training problem, identifying that the common practice in the literature of constraining the parameters of the last alternative to zero is suboptimal with respect to the so-called restricted training problem. Furthermore, the performance of KLR with the Nyström approximation was rigorously analysed, characterising the solution of Nyström KLR, and establishing the first error bound.
Secondly, the study assessed the feasibility of using the Nyström approach within KLR to analyse large datasets. This technique significantly reduced kernel matrix size, while upholding the performance of the model. In the datasets used for this comparison, performance superior to that of traditional discrete choice models was achieved, such as MNL, while reducing the size of the kernel matrices by a factor of between $110$ and $160$, compared to the original KLR model. A comprehensive comparison of four Nyström methods demonstrated their potential for efficiently handling datasets exceeding $200,000$ observations. These four methods encompass distinct sampling approaches, including a uniform, $k$-means, and two non-uniform strategies based on leverage scores. They were tested on a multi-categorical classification problem within the domain of transport mode choice, using large datasets. The $k$-means Nyström strategy emerged as particularly effective, particularly when a modest number of landmark points is used to approximate the kernel matrix. However, it was found that when the number of landmark points increases, all strategies converge to the optimal rank-$C$ approximation of Nyström.
Thirdly, the performance of Nyström-based strategies was benchmarked against traditional methods commonly applied in the transport field, such as MNL and state-of-the-art ML techniques. The results concluded that while effective, the Nyström KLR models did not surpass the superior performance of methods like XGBoost. However, a noteworthy observation is that top-performing methods like ensemble approaches (XGBoost or RF) often struggle in estimating the derivatives of probabilities, as reported by Martín-Baos et al. (2023), which is crucial for extracting econometric indicators. The KLR approach allows the derivatives of the probabilities to be estimated accurately (Martín-Baos et al., 2020), and this is what motivates further research in the direction of obtaining econometric information using these models.
@@
Lastly, datasets applied in discrete choice modelling, such as survey datasets, might exhibit imbalances due to unequal representation of alternatives in the population. The study shows that the ML methods achieve lower precision scores on the less represented classes. Addressing these imbalances is another avenue for future research. Exploring data balancing techniques and investigating their impact on model performance holds promise for further refinement. In essence, this study contributes valuable insights into the effectiveness and limitations of Nyström KLR in handling large datasets, paving the way for continued advancements in the field.
@@ Acknowledgements
This work was supported by grant PID2020-112967GB-C32 funded by MCIN/AEI/10.13039/501100011033 and by ERDF A way of making Europe. Additional funding was provided by grant 2022-GRIN-34249 from the University of Castilla-La Mancha and ERDF. We would also like to express our gratitude to the Transport and Mobility Laboratory at EPFL for their support of this work during the research stay of Martín-Baos.
@@
Acronyms
CEL
Cross-Entropy Loss
DCA
Discrete Classification Accuracy
DCM
Discrete Choice Models
DNN
Deep Neural Network
GD
Gradient Descent
GMPCA
Geometric Mean Probability of Correct Assignment
GP
Gaussian Process
KLR
Kernel Logistic Regression
ML
Machine Learning
MLE
Maximum Likelihood Estimation
MLP
Multilayer Perceptron
MNL
Multinomial Logit
NL
Nested Logit Model
NN
Neural Network
PMLE
Penalised Maximum Likelihood Estimation
RF
Random Forests
RKHS
Reproducing Kernel Hilbert Spaces
RNLL
Regularised Negative Log-Likelihood
RUM
Random Utility Model
SGD
Stochastic Gradient Descent
SVD
Singular Value Decomposition
SVM
Support Vector Machine
WTP
Willingness to Pay
XGBoost
Extreme Gradient Boosting
@@ Appendix A Analysis of the KLR model overspecification under RNLL
Applying the MLE method to KLR (by setting $\lambda=0$), it will not be possible to univocally determine the function $f_{i}(\mathbf{x})$, since there are infinite functions $f^{*}_{i}(\mathbf{x})$ leading to the same probability functions $\mathbf{p}_{\cdot i}(\boldsymbol{\alpha}_{\cdot\cdot})$ and, therefore, the log-likelihood value remains constant. This leads to the overspecification of the MLE problem, resulting in an infinite number of optimal solutions. To address this, a potential solution involves setting the parameter value for the last alternative, $I$, to $\mathbf{0}$. This choice is equivalent to setting $\boldsymbol{\beta}=-\boldsymbol{\alpha}_{I}$ yields another optimal MLE solution, where the parameter vector of the last alternative is $\mathbf{0}$. This restriction does not compromise the quality of the solution of the training problem, but the analysis is not correct when $\lambda>0$.
@@
From the fact that *$\boldsymbol{\aleph}^{*}_{\cdot\cdot}\in L(\mathbf{K})$, and from the optimality of $\boldsymbol{\aleph}^{*}_{\cdot\cdot}$ for the problem (44), Equation (45) leads to the optimality of $\boldsymbol{\aleph}^{*}_{\cdot\cdot}$ for the restricted problem $\min_{\boldsymbol{\alpha}_{\cdot\cdot}\in L(\mathbf{K})}\mathcal{L}(\boldsymbol{\alpha}_{\cdot\cdot}).$
@@
where $\mathbf{U},\boldsymbol{\Sigma}$ are an $N\times R$ column-orthogonal matrix and an $R\times R$ diagonal matrix with nonnegative entries, respectively. Assuming that $\hbox{rank}(\mathbf{K})=R$ then $\mathbf{K}$ has $\sigma_{1}>\cdots>\sigma_{R}>0$ eigenvalues and $\boldsymbol{\Sigma}=\hbox{diag}(\sigma_{1}\cdots\sigma_{R})$.
@@
${}_{\blacksquare}$
@@ Appendix B Error bounds for Nyström KLR
$\mathbf{K}\underset{\mathbf{z}_{i}}{\underbrace{\left[N\lambda\boldsymbol{\aleph}^{*}_{\cdot i}-\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})+\boldsymbol{y}_{\cdot i}\right]}}=\mathbf{0},\hskip 28.45274pti=1,\cdots,I.$ (57)
@@
$\widehat{\mathbf{K}}\left[N\lambda\boldsymbol{\aleph}^{*}_{\cdot i}-\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})+\boldsymbol{y}_{\cdot i}\right]=\mathbf{0},\hskip 28.45274pti=1,\cdots,I.$ (58)
@@
The relationship (*) is satisfied by the definition of $\mathbf{A}$. Now consider the relationship (**)(**)( * * ). Given $\widehat{\boldsymbol{\chi}}^{*}_{\cdot\cdot}\in L(\mathbf{KP})\Rightarrow\widehat{\boldsymbol{\chi}}^{*}_{\cdot\cdot}=\mathbf{KPD}$ with $\mathbf{D}\in\mathbb{R}^{C\times I}$, then:
@@
$\widehat{\mathbf{K}}\left[N\lambda\left[\widehat{\boldsymbol{\chi}}^{*}_{\cdot i}+\widehat{\boldsymbol{\chi}}^{*\perp}_{\cdot i}\right]-\widehat{\mathbf{p}}_{\cdot i}(\widehat{\boldsymbol{\chi}}^{*}_{\cdot\cdot})+\boldsymbol{y}_{\cdot i}+\boldsymbol{\varepsilon}\right]=\mathbf{0},\hskip 28.45274pti=1,\cdots I.$ (67)
@@
$\widehat{\mathbf{K}}\left[N\lambda\widehat{\boldsymbol{\chi}}^{*}_{\cdot i}-\widehat{\mathbf{p}}_{\cdot i}(\widehat{\boldsymbol{\chi}}^{*}_{\cdot\cdot})+\boldsymbol{y}_{\cdot i}\right]=\mathbf{0},\hskip 28.45274pti=1,\cdots,I.$ (68)
@@
${\mathcal{P}}^{\perp}_{L(\mathbf{K})}\left(\left[N\lambda\boldsymbol{\aleph}^{*}_{\cdot i}-\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})+\boldsymbol{y}_{\cdot i}\right]\right)={N\lambda\boldsymbol{\aleph}^{*}_{\cdot i}+\mathcal{P}}^{\perp}_{L(\mathbf{K})}\left(\left[-\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})+\boldsymbol{y}_{\cdot i}\right]\right)=\mathbf{0}$ (75)
@@
$\|{\boldsymbol{\aleph}^{*}_{\cdot i}\|_{p}=\frac{1}{N\lambda}\|\mathcal{P}}^{\perp}_{L(\mathbf{K})}\left(-\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})+\boldsymbol{y}_{\cdot i}\right)\|_{p}\leq\frac{1}{N\lambda}\|-\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})+\boldsymbol{y}_{\cdot i}\|_{p}\leq\frac{1}{N\lambda}\left[\,\|\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})\|_{p}+\|\boldsymbol{y}_{\cdot i}\|_{p}\,\right]$

### v3

@@ 6.4 Comparison of optimisation techniques for Nyström KLR
An important consideration of this study is that to obtain the graphs presented, it was necessary to report the value of the objective function at every iteration of the estimation methods. However, obtaining the value of the objective function using the KLR method with large datasets is computationally expensive. Consequently, all estimation methods display CPU times that surpass those that would have been the case if the objective function’s value were not reported, or if it were computed only at intervals of $\Delta_{t}$ iterations. Since this calculation is performed in all methods, it does not pose a problem when comparing them, which is the purpose of this experiment.
Another significant finding of this study is the consequence of Theorem 3.1, particularly Remark 3.1. The training problem of KLR is overspecified in terms of the parameters. To expedite convergence, the traditional approach is to set the parameters for one alternative to zero, i.e., $\boldsymbol{\alpha}^{*}_{\cdot\cdot}\in\boldsymbol{\Theta}=\left\{\boldsymbol{\alpha}_{\cdot\cdot}\in\mathbb{R}^{N\times I}:{\alpha}_{\cdot I}=\mathbf{0}\right\}$. However, in this study, the training problem formulation given by Equation (7) is adopted instead of the commonly used restriction $\boldsymbol{\alpha}_{\cdot\cdot}\in\boldsymbol{\Theta}$. For this reason, if the constraint ${\alpha}_{\cdot I}=\mathbf{0}$ is imposed, a suboptimal solution is obtained.
To demonstrate this, we repeated the previous experiment while enforcing ${\alpha}_{\cdot I}=\mathbf{0}$. Table 4 reports the final loss values from training, the CPU estimation time, and the GMPCA index for the test set. Comparing the loss values with those in Table 3, it becomes clear that this approach results in a suboptimal solution, as none of the optimisation methods reach the previous performance levels. Additionally, when comparing the GMPCA index on the test set with those obtained for the k-means Nyström KLR model in Table 1, we can observe that the performance on the test dataset is also inferior, indicating a reduction in the model’s generalisation capacity.
Table 4: Optimisation results with $\boldsymbol{\alpha}^{*}_{\cdot\cdot}\in\boldsymbol{\Theta}=\left\{\boldsymbol{\alpha}_{\cdot\cdot}\in\mathbb{R}^{N\times I}:{\alpha}_{\cdot I}=\mathbf{0}\right\}$
LPMC NTS
Final Loss GMPCA CPU Time (s) Final Loss GMPCA CPU Time (s)
L-BFGS-B 36,820.31 49.97 164.10 64,291.27 44.27 493.37
GD 40,040.16 47.36 1,068.42 71,838.21 40.92 2,648.15
Momentum 37,477.28 49.52 1,027.60 66,550.53 43.42 2,889.25
Adam 37,082.93 49.83 1,043.43 64,404.81 44.26 2,905.86
@@ 7 Conclusion and future work
Recent research indicates that ML models generally outperform traditional DCM in terms of predictive accuracy. However, some ML methods, such as ensemble techniques (XGBoost or RF), suffer from limited extrapolation capabilities and encounter numerical issues when calculating derivatives of the model’s probability function. As detailed in Martín-Baos et al. (2023), these problems impact the performance of econometric indicators derived from these models, including the calculation of WTP or market shares. Additionally, in Martín-Baos et al. (2021); Martín-Baos (2023), it was shown that KLR consistently return accurate choice probabilities and derivatives with respect to the attributes, providing essential information for behavioural analysis. For these reasons, despite not being the top-performing predictive technique, KLR emerges as a compelling method from a behavioural perspective.
@@
Firstly, a theoretical study was conducted to characterise the solution set of the KLR training problem, identifying that the common practice in the literature of constraining the parameters of the last alternative to zero is suboptimal with respect to the so-called restricted training problem. Furthermore, the parameters of KLR with the Nyström approximation was rigorously analysed, characterising the solution of Nyström KLR, and establishing the first error bound on learnt parameters. A pending issue for future developments is the establishment of the bounds for the performance metrics of the Nyström-KLR model as a function of the error bounds on the parameters.
Secondly, the study assessed the feasibility of using the Nyström approach within KLR to analyse large datasets. This technique significantly reduced kernel matrix size, while upholding the performance of the model. In the datasets used for this comparison, performance superior to that of traditional discrete choice models was achieved, such as MNL, while reducing the size of the kernel matrices by a factor of between $110$ and $160$, compared to the original KLR model. A comprehensive comparison of four Nyström methods demonstrated their potential for efficiently handling datasets exceeding $200,000$ observations. These four methods encompass distinct sampling approaches, including a uniform, $k$-means, and two non-uniform strategies based on leverage scores. They were tested on a multi-categorical classification problem within the domain of transport mode choice, using large datasets. The $k$-means Nyström strategy emerged as particularly effective, particularly when a modest number of landmark points is used to approximate the kernel matrix. However, it was found that when the number of landmark points increases, all strategies converge to the same trend. In Kumar et al. (2012), more specifically in Theorem 2, it is shown that standard Nyström method converge to the best rank-$C$ approximation of Nyström when the number of landmarks $C$ grows and therefore the graph shows that they all converge to the best rank $C-$ Nyström approximation.
Thirdly, the performance of Nyström-based strategies was benchmarked against traditional methods commonly applied in the transport field, such as MNL and state-of-the-art ML techniques. The results concluded that while effective, the Nyström KLR models did not surpass the superior performance of methods like XGBoost. However, as indicated previously, the top-performing methods, like ensemble approaches, often struggle in estimating the derivatives of probabilities, which is crucial for extracting econometric indicators. The KLR approach allows the derivatives of the probabilities to be estimated accurately, and this is what motivates further research in the direction of obtaining econometric information using these models.
@@
Lastly, datasets applied in discrete choice modelling, such as survey datasets, might exhibit imbalances due to unequal representation of alternatives in the population. Although not reported, the numerical tests also show that the ML methods achieve lower precision scores on the less represented classes. Addressing these imbalances is another avenue for future research. Exploring data balancing techniques and investigating their impact on model performance holds promise for further refinement. In essence, this study contributes valuable insights into the effectiveness and limitations of Nyström KLR in handling large datasets, paving the way for continued advancements in the field.
@@ Acknowledgements
This work was supported by grant PID2020-112967GB-C32 funded by MICIU/AEI/10.13039/501100011033 and by ERDF A way of making Europe. Additional funding was provided by grant 2022-GRIN-34249 from the University of Castilla-La Mancha and ERDF. We would also like to express our gratitude to the Transport and Mobility Laboratory at EPFL for their support of this work during the research stay of Martín-Baos.
@@ Appendix A Analysis of the KLR model overspecification under RNLL
Applying the MLE method to KLR (by setting $\lambda=0$), it will not be possible to univocally determine the function $f_{i}(\mathbf{x})$, since there are infinite functions $f^{*}_{i}(\mathbf{x})$ leading to the same probability functions $\mathbf{p}_{\cdot i}(\boldsymbol{\alpha}_{\cdot\cdot})$ and, therefore, the log-likelihood value remains constant. This leads to the overspecification of the MLE problem, resulting in an infinite number of optimal solutions. To address this, a potential solution involves setting the parameter value for the last alternative, $I$, to $\mathbf{0}$. This choice is equivalent to setting $\boldsymbol{\beta}=-\boldsymbol{\alpha}_{I}$ yields another optimal MLE solution, where the parameter vector of the last alternative is $\mathbf{0}$. Setting $\boldsymbol{\alpha}_{\cdot I}=\mathbf{0}$ does not compromise the quality of the solution of the training problem when $\lambda=0$ but it does affect the quality of the solution when $\lambda>0$.
@@
From the fact that $\boldsymbol{\aleph}^{*}_{\cdot\cdot}\in L(\mathbf{K})$, and from the optimality of $\boldsymbol{\aleph}^{*}_{\cdot\cdot}$ for the problem (46), Equation (47) leads to the optimality of $\boldsymbol{\aleph}^{*}_{\cdot\cdot}$ for the restricted problem $\min_{\boldsymbol{\alpha}_{\cdot\cdot}\in L(\mathbf{K})}\mathcal{L}(\boldsymbol{\alpha}_{\cdot\cdot}).$
@@
where $\mathbf{U},\boldsymbol{\Sigma}$ are an $N\times R$ column-orthogonal matrix and an $R\times R$ diagonal matrix with nonnegative entries, respectively. Assuming that $\hbox{rank}(\mathbf{K})=R$ then $\mathbf{K}$ has $\sigma_{1}\geq\cdots\geq\sigma_{R}>0$ eigenvalues and $\boldsymbol{\Sigma}=\hbox{diag}(\sigma_{1}\cdots\sigma_{R})$.
@@
Equation (15) is derived by applying matrix differentiation rules and can be found in the doctoral thesis of one of the authors (Martín-Baos, 2023). Moreover, a derivation for the case of binary classification problems can also be found in Ouyed and Allili (2018).
@@ Appendix B Error bounds for Nyström KLR
$\mathbf{K}\underset{\mathbf{z}_{i}}{\underbrace{\left[N\lambda\boldsymbol{\aleph}^{*}_{\cdot i}+\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})-\boldsymbol{y}_{\cdot i}\right]}}=\mathbf{0},\hskip 28.45274pti=1,\cdots,I.$ (59)
@@
$\widehat{\mathbf{K}}\left[N\lambda\boldsymbol{\aleph}^{*}_{\cdot i}+\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})-\boldsymbol{y}_{\cdot i}\right]=\mathbf{0},\hskip 28.45274pti=1,\cdots,I.$ (60)
@@
The relationship (*) is satisfied by the definition of $\mathbf{A}$. Now consider the relationship ()(**)( ∗ ∗ ). Given $\widehat{\boldsymbol{\chi}}^{*}_{\cdot\cdot}\in L(\mathbf{KP})\Rightarrow\widehat{\boldsymbol{\chi}}^{*}_{\cdot\cdot}=\mathbf{KPD}$ with $\mathbf{D}\in\mathbb{R}^{C\times I}$, then:
@@
$\widehat{\mathbf{K}}\left[N\lambda\left[\widehat{\boldsymbol{\chi}}^{*}_{\cdot i}+\widehat{\boldsymbol{\chi}}^{*\perp}_{\cdot i}\right]+\widehat{\mathbf{p}}_{\cdot i}(\widehat{\boldsymbol{\chi}}^{*}_{\cdot\cdot})-\boldsymbol{y}_{\cdot i}+\boldsymbol{\varepsilon}\right]=\mathbf{0},\hskip 28.45274pti=1,\cdots I.$ (69)
@@
$\widehat{\mathbf{K}}\left[N\lambda\widehat{\boldsymbol{\chi}}^{*}_{\cdot i}+\widehat{\mathbf{p}}_{\cdot i}(\widehat{\boldsymbol{\chi}}^{*}_{\cdot\cdot})-\boldsymbol{y}_{\cdot i}\right]=\mathbf{0},\hskip 28.45274pti=1,\cdots,I.$ (70)
@@
${\mathcal{P}}^{\perp}_{L(\mathbf{K})}\left(\left[N\lambda\boldsymbol{\aleph}^{*}_{\cdot i}+\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})-\boldsymbol{y}_{\cdot i}\right]\right)={N\lambda\boldsymbol{\aleph}^{*}_{\cdot i}+\mathcal{P}}^{\perp}_{L(\mathbf{K})}\left(\left[\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})-\boldsymbol{y}_{\cdot i}\right]\right)=\mathbf{0}$ (77)
@@
$\|{\boldsymbol{\aleph}^{*}_{\cdot i}\|_{p}=\frac{1}{N\lambda}\|\mathcal{P}}^{\perp}_{L(\mathbf{K})}\left(\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})-\boldsymbol{y}_{\cdot i}\right)\|_{p}\leq\frac{1}{N\lambda}\|\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})-\boldsymbol{y}_{\cdot i}\|_{p}\leq\frac{1}{N\lambda}\left[\,\|\mathbf{p}_{\cdot i}(\boldsymbol{\aleph}^{*}_{\cdot\cdot})\|_{p}+\|\boldsymbol{y}_{\cdot i}\|_{p}\,\right]$